# Physically-based Lighting Augmentation for Robotic Manipulation

### v1

@@ 4.1 Visual Lighting Quality Evaluation
We apply lighting augmentation to 10 randomly selected sequences (600 frames) from the real-world human demonstrations in Sec.4.2 using our method and IC-Light [68], a widely recognized text-prompt-based relighting method. Ground-truth episodes are recorded under a side-mounted blue LED panel light under the same task execution trajectories. Two sets of metrics are adopted: (1) LPIPS [69] and SSIM [70] are computed between the ground-truth and relit image sequences to assess structural consistency and fidelity. (2) Temporal LPIPS and Temporal SSIM are computed between consecutive frames in the relit sequences to assess temporal consistency. Since IC-Light does not support video relighting, we apply it frame-by-frame using a fixed random seed and the prompt “blue LED light” for consistency. We perform relighting using an environment map approximated from the ground truth episodes with [18]. Qualitative and quantitative evaluation results can be found in Fig. 4. Our method achieves better performance by estimating scene-level properties, which are more compatible with robotic settings, whereas IC-Light’s foreground/background separation is better suited for human relighting. Our temporal propagation module further improves visual consistency by reducing abrupt changes between frames.
@@
Figure 4: Qualitative (top) and quantitative (bottom) evaluation of lighting augmentation. Although exceeding the baseline, the lighting in our augmented episodes appears more diffused compared to the ground truth, as indicated by weaker side shadows. This discrepancy mainly results from the rough environment map approximation (32×1632$\times$ 1632 × 16 resolution) without specialized graphics equipment.
4.2 Real-world Evaluation
Lighting Setup.
White light with minimal shadows (Ambient, Fig. 5(a)) is used during expert human demonstration collection for policy training. During evaluation, six lighting conditions are designed using a programmable RGB LED panel. This includes two sets with different evaluation purposes: (1) Daily-use conditions — Side (strong shadows), Regional (blue regional highlight), and Dim (low-light); (2) Artificial RGB lighting — Red, Green, and Blue — with illuminance levels of 4,600 lux and RGB values (255, 0, 0), (0, 255, 0), and (0, 0, 255), respectively, designed to evaluate each policy under single-source lighting. These colors are chosen based on the principle of light transport linearity, which allows any colored illumination to be reconstructed as their linear combination. Examples are shown in Fig. 5(b).
Manipulation Tasks.
Pear: Grasp a randomly placed green plastic pear using a vacuum gripper and place it on a plate. Square: Grasp a randomly placed black square nut by its handle using a parallel jaw gripper and insert it into a matching red square peg. To better assess performance under challenging lighting conditions, we report performance separately for the reaching (R) subtask, which involves reaching the area of the pear or square nut, and the pick-and-place (PnP) outcome, defined as successfully picking up the object and placing it in the corresponding plate or peg.
@@ 4.2 Real-world Evaluation
Figure 5: (a) Experimental setup under Ambient lighting. (b) The six lighting conditions used for generalization gap evaluation. (c) Objects used in the two manipulation tasks.
Implementation Details.
We perform behavior cloning using ResNet18 [71] and Multi-Layer Perceptron (BC-MLP) to map features to actions. Random cropping (Crop) is used as the default augmentation due to its widespread use and proven effectiveness [5, 39]. We evaluate two variants: +Jitter, which adds color jitter, and +Ours, which adds our proposed lighting augmentation. Color jitter is selected as a baseline due to its ability to mitigate the impact of color and illumination variation. Training and evaluation are conducted on a 7-DoF UFactory XArm7 robot observed by a single externally mounted RGB camera. Each task uses 200 human demonstrations to train the baseline policies Crop and +Jitter. For +Ours, 10 demonstrations are randomly selected and augmented with six environment maps approximating the six evaluation lighting conditions, producing 260 training episodes per task for policy training. See the supplementary material for full implementation details.
Evaluation Protocol.
For each combination of task $\tau$, lighting condition L, and augmentation method Aug, we conduct 20 real-world evaluations. The success rate for each combination is denoted as $S_{\textit{L}}^{\textit{Aug}}(\tau)$. We use BC-MLP trained and evaluated under Ambient lighting with Crop augmentation as the base policy. The corresponding subtask success rates are: $S_{\textit{Ambient}}^{\textit{Crop}}(\textit{Pear}\text{-}\textit{R})$ = 1.00, $S_{\textit{Ambient}}^{\textit{Crop}}(\textit{Pear}\text{-}\textit{PnP})$ = 0.60, $S_{\textit{Ambient}}^{\textit{Crop}}(\textit{Square}\text{-}\textit{R})$ = 1.00, $S_{\textit{Ambient}}^{\textit{Crop}}(\textit{Square}\text{-}\textit{PnP})$ = 0.85. Performance degradation under unseen evaluation lighting conditions is measured by the Generalization Gap, defined as the normalized drop in success rate relative to the base policy:
$\textit{Generalization Gap}=(S_{\textit{Ambient}}^{\textit{Crop}}(\tau)-S_{\textit{L}}^{\textit{Aug}}(\tau))/S_{\textit{Ambient}}^{\textit{Crop}}(\tau).$ (7)
Results.
Tab. 1 presents the generalization gap of BC-MLP under six unseen lighting conditions across 720 real-world evaluations. Averaged over both subtasks across the two tasks, our method outperforms Crop by 40.1% and +Jitter by 35.7%. In the daily-use lighting set, Crop shows a major drop in performance, except under Dim. While +Jitter performs reasonably on the reaching subtask, it consistently fails in the final pick-and-place outcome. +Ours improves performance across both subtasks under these conditions. In the artificial RGB lighting set, both Crop and +Jitter fail in most cases, even on the reaching subtask. Interestingly, +Jitter displays strong preferences for different RGB lighting in the two tasks, likely due to its use of global color shifts. Our method maintains strong reaching performance but struggles with pick-and-place in these more extreme lighting conditions. Common failure modes for Crop and +Jitter involve moving directly to the plate or peg without interacting with the pear or nut. Failures in +Ours often involve reaching the pear or nut with the gripper open and hovering just above the object, without executing a pick action. This may relate to the altered object appearance under high-intensity lighting (4,600 lux), where high reflectivity could disrupt perception. See the limitations section for further details and the supplementary video for examples of common failure modes across different methods.
@@ Results.
Table 1: Generalization Gap ($\downarrow$) with different augmentation methods under unseen lighting. The best-performing policy is in bold, and the second-best is underscored.
Lighting Crop +Jitter +Ours
Pear Square Pear Square Pear Square
R PnP R PnP R PnP R PnP R PnP R PnP
Side 1.00 1.00 1.00 1.00 0.70 1.00 0.85 1.00 0.92 1.00 0.10 0.53
Regional 0.95 1.00 1.00 1.00 0.55 1.00 0.50 1.00 0 0.67 0 0.47
Dim 0 0.16 0.05 0.35 0.1 0.83 0.05 0.47 0 0.34 0 0.06
Red 1.00 1.00 1.00 1.00 0.70 1.00 0.90 1.00 0.40 1.00 0.10 1.00
Green 0.85 1.00 0.90 1.00 1.00 1.00 0.95 1.00 0 1.00 0 0.88
Blue 1.00 1.00 0.95 1.00 1.00 1.00 0.55 1.00 0.15 0.92 0.05 1.00
4.3 Ablation Study
Figure 6: Ablation study: (a) effect of the number of lighting conditions; (b) effect of the number of augmented episodes.
Impact of the Number of Lighting Conditions for Augmentation.
We perform augmentation using 1 (Red), 3 (Red, Green, Blue), 6 (all conditions from Sec. 4.2), and 9 lighting conditions (6 plus three additional lights with RGB values (255, 255, 0), (255, 0, 255), and (0, 255, 255)). Evaluation is performed in single-source Red light to control for lighting composition effects and isolate augmentation impact. Reaching success rates on Square task, averaged over 10 trials, are shown in Fig. 6(a). Including the currently observed lighting (Red) in augmentation immediately boosts performance (from 0%). However, performance degrades when too many unrelated lighting conditions are added, likely due to the reduced presence of target augmentations (Red) in the expanded set. See Sec. 7 for further discussion.
Impact of the Number of Episodes for Augmentation.
@@ 4.3 Ablation Study
We select 10, 25, 75, and 100 episodes from 200 demonstrations and augment each with six lighting conditions. Evaluation follows the same setup as above. As shown in Fig. 6(b), performance is relatively unaffected by the number of augmented episodes, with no signs of overfitting — possibly due to the lighting diversity within each episode. To save computation, we use the minimal setting of 10 episodes for augmentation.
5 Applications
@@
Through scene decomposition via inverse rendering, we extend our framework to augment three additional environmental factors (background, object texture, and distractors) beyond lighting (Fig. 7(a)) on existing datasets, as demonstrated on BridgeData v2 [72].
5.1 Background Augmentation
@@
Background appearance is tied to global illumination; therefore, we render segmented scene geometry with different environment maps to create new backgrounds with corresponding lighting. Scene geometry is reconstructed by triangulating the depth map $D^{*}$ from Sec. 3.2 into a mesh $\Lambda$:
@@
[…]
Figure 7: Examples of applications: (a) Lighting augmentation with different environment maps; (b) Background augmentation with different environment maps; (c) Texture augmentation via albedo adjustment; (d) Gravity-based distractor placement, with distractor mesh sourced from [74].
5.2 Object Texture Augmentation
@@
Robotic tasks like grasping and pushing rely on consistent object geometry and frictional properties  [75, 76]. We address this by applying object texture augmentation with preserved visual roughness and geometry. Following the mesh segmentation in Sec. 5.1, we segment the object mesh and render with altered material properties. Specifically, albedo $A^{*}$ sets the base color, roughness $R^{*}$ controls surface scattering, and metallic $M^{*}$ defines the metallic effect. By adjusting only the albedo $A^{*}$ while keeping roughness $R^{*}$ and metallic $M^{*}$ fixed, and reapplying Eq. 4, we achieve object texture augmentation with consistent visual roughness and geometry. Examples are shown in Fig. 7(c).
5.3 Distractor Placement in Cluttered Environments
@@
Previous methods typically place distractors via inpainting using text-driven generative models [17, 16, 4]. We take a different approach by leveraging the meshes of the scene and distractors for physically plausible placement in Blender [77]. As shown in Fig. 7(d), we first align the surface normal with Blender’s default gravity axis (Z-axis). This requires the user to manually select a mesh face whose surface normal is parallel to the gravity direction using the Blender API. A random position within the bounding box of $\Lambda$ (Eq. 8) is sampled, and the distractor mesh — assigned a pseudo-mass of 1kg at its geometric center — is dropped to simulate its motion under gravity. The final placement of the distractor is determined by its stable resting state following the gravity-based simulation. The full simulation process is available in the supplementary video.
6 Conclusion
We propose the first framework that applies physically-based lighting augmentation on real-world human demonstrations for robotic manipulation. By decomposing robotic scenes into geometric and material properties, we perform lighting augmentation on a single frame and propagate it across the entire demonstration using finetuned Stable Video Diffusion. We validate our framework through qualitative and quantitative visual quality evaluations, 720 real-world trials under six varied lighting conditions using a 7-DoF robot, and complementary ablation studies. We further showcase augmentations on three additional environmental factors enabled by our framework. This work marks an initial step toward integrating physical realism into visual augmentation, with the potential to assist future research on improving imitation learning generalization under environmental variation.
7 Limitations and Future work
Environment Lighting Approximation.
@@ 7 Limitations and Future work
Although we provide physically-based lighting augmentation by simulating light-material interactions, the absence of an environment map that perfectly reconstructs the currently observed lighting leads to reduced performance. One example is shown in Fig. 8, where under high-intensity red LED lighting (4600 lux), the plastic pear and plate exhibit strong reflections, differing from our augmentation. The discrepancy results from the side-mounted high-intensity LED producing strong directional lighting, while the approximated environment map failed to capture this and provides primarily diffuse illumination with a shifted color tone. This is partially due to the limited precision of the single-view-based environment map approximation (32×1632$\times$ 1632 × 16 resolution). Capturing current lighting using graphical equipment [62] to generate accurate environment maps could mitigate this issue. Future work will focus on improving lighting approximation and deriving high-precision environment maps from single-view input for robotic scenes.
@@
Efficiency.
The current temporal propagation module only applies to lighting augmentation and does not extend to the other three environmental factors discussed in Sec. 5, as they are not captured by the curated SVD fine-tuning dataset. Consequently, augmentation for the additional factors is currently performed frame by frame, which makes the proposed applications more suitable for offline augmentation of human demonstration data. Future work will focus on improving efficiency through selective optimization of geometric and material properties, taking into account environmental similarities in robotic data collection settings.
A Universal Lighting Composition.
This paper focuses on using the proposed lighting augmentation framework to improve performance under currently observed lighting conditions, as the specific and quantified effects of different lighting on robot learning remain largely unknown. Future work will aim to systematically study how various lighting conditions impact policy performance and, based on this understanding, develop effective augmentation strategies for constructing a universal lighting composition to support lighting-invariant policies.
Visual-Physical Gap.
The term physically-based in this paper refers to physically-based rendering. While we follow physically accurate light–material interactions from a computer graphics perspective, these effects are purely visual and do not affect motion dynamics, which is a critical aspect of robotic manipulation.

### v2

@@ IV-A Visual Lighting Quality Evaluation
We apply novel lighting generation to 10 randomly selected sequences (600 frames) from the real-world human demonstrations in Sec.IV-B using our method and IC-Light [42], a widely recognized text-prompt-based relighting method. Ground-truth episodes are recorded under a side-mounted blue LED panel light under the same task execution trajectories. Two sets of metrics are adopted: (1) LPIPS [43] and SSIM [44] are computed between the ground-truth and relit image sequences to assess structural consistency and fidelity. (2) Temporal LPIPS and Temporal SSIM are computed between consecutive frames in the relit sequences to assess temporal consistency. Since IC-Light does not support video relighting, we apply it frame-by-frame using a fixed random seed and the prompt “blue LED light” for consistency. We perform relighting using an environment map approximated from the ground truth episodes with [12]. Qualitative and quantitative evaluation results can be found in Fig. 5. Our method performs better by preserving structural similarity with the original frames. The use of environment maps for target lighting provides precise control over relighting, making it more effective than textual prompts when approximating real-world data under varying lighting conditions. The temporal propagation module further enhances visual consistency by reducing abrupt changes between frames. Beyond visual quality evaluation, IC-Light is also included in subsequent real-world comparisons in Sec. IV-B.
@@
Figure 5: Qualitative (top) and quantitative (bottom) evaluation of lighting generation. Although exceeding the baseline, the lighting in our generated episodes appears more diffused compared to the ground truth, as indicated by weaker side shadows. This discrepancy mainly results from the rough environment map approximation ($32\times 16$ resolution) without specialized graphics equipment.
@@
Figure 6: (a) Experimental setup under Original lighting. (b) The six lighting conditions used for policy evaluation under unseen lighting. (c) Objects used in the two manipulation tasks.
@@
TABLE I: Success rate ($\uparrow$) with different generation methods under unseen lighting. The best-performing policy is in bold, and the second-best is underscored.
Lighting Crop Jitter IC-Light Ours
Pear Square Pear Square Pear Square Pear Square
R PnP R PnP R PnP R PnP R PnP R PnP R PnP R PnP
Original 1 0.60 1 0.85 / / / / / / / / / / / /
Side 0 0 0 0 0.3 0 0.15 0 0.05 0 0.2 0 0.05 0 0.9 0.4
Regional 0.05 0 0 0 0.45 0 0.5 0 0.10 0 0.15 0 1.0 0.2 1.0 0.45
Dim 1.0 0.5 0.95 0.55 0.9 0.1 0.95 0.45 0.20 0 0.25 0 1.0 0.4 1.0 0.8
Red 0 0 0 0 0.3 0 0.1 0 0.45 0 0.3 0 0.6 0 0.9 0
Green 0.15 0 0.1 0 0 0 0.05 0 0.25 0 0.1 0 1.0 0 1.0 0.1
Blue 0 0 0.05 0 0 0 0.45 0 0.4 0 0.15 0 0.85 0.05 0.95 0
  • The vanilla BC-MLP with random cropping augmentation (Crop) evaluated under Original lighting is added as the reference of the task difficulty.
IV-B Real-world Evaluation
This experiment examines whether imitation learning policies perform better under unseen lighting when trained on episodes that approximate such conditions.
IV-B1 Implementation Details
Lighting Setup
White light (Original, Fig. 6(a)) used in most existing robotic manipulation datasets is adopted during expert human demonstration collection for policy training. During evaluation, six lighting conditions are designed using a programmable RGB LED panel. This includes two sets with different evaluation purposes: (1) Daily-use conditions—Side (strong shadows), Regional (blue regional highlight), and Dim (low-light); (2) Artificial RGB lighting—Red, Green, and Blue—with illuminance levels of 4,600 lux and RGB values (255, 0, 0), (0, 255, 0), and (0, 0, 255), respectively, designed to evaluate each policy under single-source lighting. These colors are chosen based on the principle of light transport linearity, which allows any colored illumination to be reconstructed as their linear combination. Examples are shown in Fig. 6(b).
Manipulation Tasks
(1) Pear, grasp a randomly placed plastic pear using a vacuum gripper and place it on a plate. (2) Square, grasp a randomly placed black square nut by its handle using a parallel jaw gripper and insert it into a red square peg. To better assess performance under challenging lighting conditions, we report performance separately for the reaching (R) subtask, which involves reaching the area of the pear or square nut, and the pick-and-place (PnP) outcome, defined as successfully picking up the object and placing it in the corresponding plate or peg.
Baselines
We perform behavior cloning with ResNet18 [45] and a Multi-Layer Perceptron (BC-MLP) to map features to actions. This architecture is chosen for its widespread use [46, 47] and simplicity, providing a clean testbed to evaluate the impact of generated data. Training is conducted on a 7-DoF UFactory XArm7 robot observed by a single externally mounted RGB camera. We compare four baselines: (a) Crop, vanilla BC-MLP with random cropping, trained on 200 human demonstration episodes under Original lighting; (b) Jitter, random cropping plus color jitter, trained on the same 200 episodes; (c) IC-Light, trained on the original 200 episodes plus generated episodes approximating unseen lighting conditions with IC-Light; and (d) Ours, trained on the original 200 episodes plus generated episodes from our method. For IC-Light and Ours, 10 demonstrations are randomly selected and relit with six environment maps matching the six evaluation conditions (derived through inverse rendering), yielding 260 training episodes per task. Color jitter is included as a baseline due to its ability to mitigate color and illumination variation.
Evaluation Protocol
For each combination of task, lighting condition, and method, we conduct 20 real-world evaluations and report their success rates. Pear (Fig. 1(a)) and Square (Fig. 6(a)) differ slightly in background, and the wrist camera is included only for automatic evaluation, not for training purposes. All baselines within a task are evaluated in the same environment.
IV-B2 Results
Tab. I presents the success rates of BC-MLP under six unseen lighting conditions across 1,000 real-world evaluations. Averaged over both subtasks across the two tasks, our method outperforms Crop by 38.75%, Jitter by 33.13% and IC-Light by 41.88%. In the daily-use lighting set, Crop shows a major drop in performance, except under Dim. While Jitter performs reasonably on the reaching subtask, it consistently fails in the final pick-and-place outcome. IC-Light exhibits reduced performance on daily-use lighting but better performance on RGB lighting, possibly because it generates episodes that align more closely with RGB lighting. Ours improves performance across both subtasks under these conditions. In the artificial RGB lighting set, both Crop and Jitter fail in most cases, even on the reaching subtask. Interestingly, Jitter displays strong preferences for different RGB lighting in the two tasks, likely due to its use of global color shifts. Our method maintains strong reaching performance but struggles with pick-and-place in these more extreme lighting conditions. Common failure modes for Crop, Jitter and IC-Light involve moving directly to the plate or peg without interacting with the pear or nut. Failures in Ours often involve reaching the pear or nut with the gripper open and hovering just above the object, without executing a pick action. This may relate to the altered object appearance under high-intensity lighting (4,600 lux), where high reflectivity could disrupt perception. See the limitations section for further details and the supplementary video for examples of common failure modes across different methods.
IV-C Ablation Study
We provide ablation study on individual modules of the proposed framework: Inverse Rendering (Sec. IV-C1), Temporal Propagation (Sec. IV-C2, Sec. IV-C3) and Behavior Cloning (Sec. IV-C4).
IV-C1 Material Property Ablation
As shown in Tab. II, we ablate material properties by individually masking estimated albedo, roughness, and metallic to 0.5 (originally in [0, 1]) and evaluate the trained policy’s success rate (R/PnP) over 10 trials on task Square under Dim. Policy trained on the original estimates is included for comparison. Results suggest albedo and roughness are critical for policy performance, while metallic has minimal effect, likely due to the non-metallic object. Masking any property leads to reduced performance.
TABLE II: Material property ablation.
Original Albedo Roughness Metallic
1.0/0.8 0.8/0 1.0/0 1.0/0.4
IV-C2 SVD Fine-tuning Data Ablation
The fine-tuning data for SVD consist of two sources: synthetic robot execution videos, which provide varied lighting across tasks ($\mathcal{D}_{1}$), and real-world execution videos, which include visual degradation to improve radiometric accuracy ($\mathcal{D}_{2}$). In Fig. 7, we show an example frame generated by SVD when fine-tuned only on $\mathcal{D}_{1}$. The result exhibits low radiometric accuracy and a noticeably cartoonish appearance.
Figure 7: SVD fine-tuning data ablation.
IV-C3 Temporal Propagation Ablation
Tab. III shows the comparison of relighting time: IC-Light, frame-wise inverse rendering (Ours (w/o SVD)), and the proposed first frame inverse rendering + SVD (Ours (w/ SVD)). With SVD, after a one-time 10-minute inverse rendering, relighting each episode takes only 20 seconds. The efficiency gain becomes greater when relighting an episode with multiple lighting, e.g. 6 relightings take just 12 minutes: $10+(20/60)\times 6$.
TABLE III: Time efficiency comparison.
Methods 1 Frame 1 Episode (61 Frames) 6 Episodes
IC-Light 10 s 10 min 1 hour
Ours (w/o SVD) 10 min 10 hours 60 hours
Ours (w/ SVD) 10 min 10 min + 20 s 12 min
IV-C4 Impact of the Number of Generated Episodes for Policy Training
@@
We select 10, 25, 75, and 100 episodes from 200 demonstrations and generate each with six lighting conditions. Evaluation is performed in single-source Red light to isolate generation impact. Reaching success rates on Square task, averaged over 10 trials. As shown in Tab. IV, performance is relatively unaffected by the number of generated episodes, with no signs of overfitting. To save computation, we use the minimal setting of 10 generated episodes for policy training.
TABLE IV: Impact of the number of generated episodes.
Num. Episodes 10 25 75 100
Success Rate (R) 0.9 0.8 0.8 1.0
V Applications
@@
Through scene decomposition via inverse rendering, we extend our framework to generate three additional environment factors (background, object texture, and distractors) beyond lighting (Fig. 8) on existing datasets, as demonstrated on BridgeData v2 [5].
Figure 8: Examples of lighting generation with different environment maps on existing open-source dataset.
V-A Background Generation
@@
Background appearance is tied to illumination; therefore, we render segmented scene geometry with different environment maps to create new backgrounds with corresponding lighting. Scene geometry is reconstructed by triangulating the depth map $D^{*}$ from Sec. III-B into a mesh $\Lambda$:
@@
where $K$ is a fixed intrinsic matrix of a pinhole camera, and $D^{*}(x,y)$ is the depth at pixel $(x,y)$. We then project 2D segmentation masks [48] onto the mesh $\Lambda$ from the default viewpoint to isolate and segment the robot arm and task area. The segmented mesh is then rendered with various environment maps to produce new backgrounds and lighting. Examples are shown in Fig.9.
Figure 9: Examples of background generation with different environment maps from [38].
V-B Object Texture Generation
@@
Robotic tasks like grasping and pushing rely on consistent object geometry and frictional properties  [49]. We address this by applying object texture generation with preserved visual roughness and geometry. Following the mesh segmentation in Sec. V-A, we segment the object mesh and render with altered material properties. Specifically, albedo $A^{*}$ sets the base color, roughness $R^{*}$ controls surface scattering, and metallic $M^{*}$ defines the metallic effect. By adjusting only the albedo $A^{*}$ while keeping roughness $R^{*}$ and metallic $M^{*}$ fixed, and reapplying Eq. 4, we achieve object texture generation with consistent visual roughness and geometry. Examples are shown in Fig. 10.
Figure 10: Examples of texture generation via albedo adjustment.
V-C Distractor Placement in Cluttered Environments
@@
Through the meshes of the scene and distractors, we conduct physically plausible distractor placement in Blender [50]. As shown in Fig. 11, we first align the surface normal with Blender’s default gravity axis (Z-axis). This requires the user to manually select a mesh face whose surface normal is parallel to the gravity direction using the Blender API. A random position within the bounding box of $\Lambda$ (Eq. 7) is sampled, and the distractor mesh (assigned a pseudo-mass of 1kg at its geometric center) is dropped to simulate its motion under gravity. The final placement of the distractor is determined by its stable resting state following the gravity-based simulation. The full simulation process is available in the supplementary video.
Figure 11: Examples of gravity-based distractor placement, with distractor mesh sourced from [51].
VI Limitations and Conclusion
VI-A Limitations and Future Work
@@
Although we provide physically-based lighting generation by simulating light–material interactions, the absence of an environment map that perfectly reconstructs the observed lighting reduces performance. An example is shown in Fig. 12, where under high-intensity red LED lighting (4600 lux), the plastic pear and plate exhibit strong reflections not captured by our generation. The discrepancy arises from the side-mounted LED producing strong directional lighting, while the approximated environment map ($32\times 16$ resolution) fails to capture this and instead provides mostly diffuse illumination with a shifted color tone. Capturing current lighting with graphical equipment [39] to generate accurate environment maps could mitigate this issue. Apart from above, our paper focuses on improving policy performance under unseen lighting by generating data that approximates such conditions, rather than developing lighting-invariant policies that function across all diverse lighting. Future work will aim to systematically study how different lighting conditions affect policy performance, combined with lighting generation techniques, toward developing strategies that enable lighting-invariant policies.
@@
VI-B Conclusion
In this paper, we propose the first framework that leverages physically-based inverse rendering for novel lighting generation on existing real-world human demonstrations. By decomposing robotic scenes into geometric and material properties, we perform lighting generation on a single frame and propagate it across the entire demonstration using finetuned Stable Video Diffusion. We validate our framework through qualitative and quantitative visual quality evaluations, 1,000 real-world trials under six varied lighting conditions using a 7-DoF robot, and ablation studies on individual modules of the proposed framework. We further showcase generations on three additional visual variations enabled by our framework. The code for our framework will be publicly released to support future research in physically-based data generation for robotic manipulation.